OpenStudy (anonymous):

solve by the substitution method 6x-2y=-50 6x+55=y

OpenStudy (anonymous):

6x-2(6x+55)=-50

OpenStudy (anonymous):

(-2*6x)+(2*55)=-12x+110 6x-12x+110=-50

OpenStudy (anonymous):

ty